# A Finite-Time Analysis of Distributed Q-Learning

### v1

@@ 4 Error analysis of distributed Q-learning : Markovian observation model
The following number of samples are required for $\mathbb{E}\left[\left\|\bar{{\bm{Q}}}_{k}-\bm{1}_{N}\otimes{\bm{Q}}^{*}\right\|_{\infty}\right]\leq\epsilon$:
$\tilde{{\mathcal{O}}}\left(\min\left\{\frac{\ln^{2}\left(\frac{1}{{\epsilon}^{2}}\right)}{{\epsilon}^{2}}\frac{t_{\mathrm{mix}}d_{\max}^{2}}{(1-\gamma)^{6}d_{\min}^{4}},\frac{\ln\left(\frac{1}{{\epsilon}}\right)}{{\epsilon}}\frac{d_{\max}\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\gamma)^{4}d_{\min}^{3}(1-\sigma_{2}({\bm{W}}))}\right\}\right).$
The proof is given in Appendix Section G.3. As in the result of i.i.d. case in Corollary 3.8, we have the dependency on $\frac{1}{1-\gamma},\frac{1}{d_{\min}}$, and $\frac{1}{1-\sigma_{2}({\bm{W}})}$ with additional factor on mixing time. The known tight sample complexity result in the single-agent case is $\tilde{{\mathcal{O}}}\left(\frac{1}{(1-\gamma)^{4}d_{\min}{\epsilon}^{2}}+\frac{t_{\mathrm{mix}}}{(1-\gamma)d_{\min}}\right)$ by Li et al. [2024], and our result leaves room for improvement. Assuming a uniform sampling scheme, i.e., $d_{\min}=d_{\max}=\frac{1}{|{\mathcal{S}}||{\mathcal{A}}|}$, and $|{\mathcal{A}}_{i}|=A$ for all $i\in[N]$ and $A\geq 2$, the sample complexity becomes $\tilde{{\mathcal{O}}}\left(\min\left\{\frac{t_{\text{mix}}}{{\epsilon}^{2}}\frac{|{\mathcal{S}}|^{2}A^{2N}}{(1-\gamma)^{6}},\frac{1}{{\epsilon}}\frac{|{\mathcal{S}}|^{\frac{5}{2}}A^{\frac{5N}{2}}}{(1-\gamma)^{4}(1-\sigma_{2}({\bm{W}}))}\right\}\right)$. We note that the exponential scaling in the action space is inevitable in the tabular setting unless we assume a particular structure on the problem [Gu et al., 2024] or consider a near-optimal solution [Qu et al., 2022].
@@ 5 Discussion
Q-function Assumption Sample complexity induced by graph structure Step-size Ours Tabular $\tilde{{\mathcal{O}}}\left(\frac{1}{1-\sigma_{2}({\bm{W}})}\frac{1}{{\epsilon}}\right)$ $\alpha\leq\min\left\{\min_{i\in[N]}[{\bm{W}}]_{ii},\frac{1}{2\tau}\right\}$  Heredia et al. [2020] LFA (15) $\tilde{{\mathcal{O}}}\left(\frac{1}{1-\sigma_{2}({\bm{W}})}\frac{1}{{\epsilon}}\right)$ $\frac{1}{(d_{\min}-\gamma^{2}d_{\max})(k+1)},\;k\in{\mathbb{N}}$  Zeng et al. [2022b] LFA (16) $\tilde{{\mathcal{O}}}\left(\frac{1}{(1-\sigma_{2}({\bm{W}}))^{2}}\frac{1}{{\epsilon}^{2}}\right)$ $\alpha=\text{poly}\left(\frac{1}{N},\frac{1}{1-\gamma},d_{\min},\tau,\sigma_{2}({\bm{W}})\right)$
Table 1: LFA stands for linear function approximation. The third column shows the dependency of the sample complexity on the graph structure. The last column shows the step-size for the convergence in expectation to hold. $\text{poly}(\cdot)$ stands for a function polynomially dependent on the input arguments.
@@
for some $\kappa>0$. We have provided examples where the above conditions in (15) and (16) are not met even in the tabular case in Appendix Section H. Furthermore, as can be seen in Table 1, our dependency on the property of the graph structure, $\frac{1}{1-\sigma_{2}({\bm{W}})}$ is comparable or better than other works, and assumes mild condition on the step-size, as discussed in Section 3.
@@ 6 Conclusion
In this paper, we have studied distributed version of Q-learning algorithm. We provided sample complexity result of $\tilde{\mathcal{O}}\left(\min\left\{\frac{1}{\epsilon^{2}}\frac{1}{(1-\gamma)^{6}d_{\min}^{4}},\frac{1}{\epsilon}\frac{\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\sigma_{2}(\bm{W}))(1-\gamma)^{4}d_{\min}^{3}}\right\}\right)$, which appears to be the first non-asymptotic result for tabular Q-learning. Future work would include improving the dependency on $\frac{1}{1-\gamma}$ and $d_{\min}$ to match the known tightest sample complexity bound of single-agent Q-learning [Li et al., 2020], and extending the result to neural network approximation scheme.
Appendix A Appendix : Notations
$\mathbb{R}^{n}$: set of real-valued $n$-dimensional vectors; $\mathbb{R}^{n\times m}$ : set of real-valued $n\times m$-dimensional matrices; $\Delta^{n}$ for $n\in{\mathbb{N}}$ : a probability simplex in $\mathbb{R}^{n}$; $[n]$ for $n\in{\mathbb{N}}$ : $\{1,2,\dots,n\}$; $\bm{1}_{n}$ : $n$-dimensional vector whose elements are all one; $\bm{0}$ : a vector whose elements are all zero with appropriate dimension; $[{\bm{A}}]_{ij}$ : $i$-th row and $j$-th column for any matrix ${\bm{A}}$; ${\bm{e}}_{j}$ : basis vector (with appropriate dimension) whose $j$-th element is one and others are all zero; $|{\mathcal{S}}|$ : cardinality of any finite set ${\mathcal{S}}$; $\otimes$ : Kronecker product between two matrices; ${\bm{a}}\geq{\bm{b}}$ for ${\bm{a}},{\bm{b}}\in\mathbb{R}^{n}$ : $[{\bm{a}}]_{i}\geq[{\bm{b}}]_{i}$ for all $i\in[n]$.
@@ Appendix D Appendix : Omitted Proofs
To prove the inequality in (17), we first prove that $1$ is the unique largest eigenvalue of ${\bm{W}}$. Noting that $\bm{1}_{N}$ is an eigenvector of ${\bm{W}}$ with eigenvalue of $1$, and $\rho({\bm{W}})\leq||{\bm{W}}||_{\infty}=1$ where $\rho(\cdot)$ is the spectral radius of a matrix, the largest eigenvalue of ${\bm{W}}$ should be one. This implies that $\sigma_{2}({\bm{W}})<1$. The multiplicity of the eigenvalue $1$ is one, which follows from the fact that ${\bm{W}}^{k}$ is a non-negative and irreducible matrix and that the largest eigenvalue of a non-negative and irreducible matrix is unique Pillai et al. [2005] from Perron-Frobenius theorem. Note that ${\bm{W}}^{k}$ is a non-negative and irreducible matrix due to the fact that the graph ${\mathcal{G}}$ is connected.
@@ E.2 Construction of upper comparison system
$\mathbb{E}\left[{\bm{S}}_{0}\right]=$ $\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}({\bm{e}}_{s_{0}}\otimes{\bm{e}}_{{\bm{a}}_{0}})(r^{i}_{1}+{\bm{e}}_{s^{\prime}_{0}}^{\top}\gamma{\bm{\Pi}}^{{\bm{Q}}^{i}_{0}}{\bm{Q}}^{i}_{0}-({\bm{e}}_{s_{0}}\otimes{\bm{e}}_{{\bm{a}}_{0}})^{\top}{\bm{Q}}^{i}_{0})\right]-\mathbb{E}\left[\frac{1}{N}\sum^{N}_{i=1}{\bm{D}}({\bm{R}}^{i}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}^{i}_{0}}{\bm{Q}}^{i}_{k}-{\bm{Q}}^{i}_{0})\right]$
@@
$\leq$ $(1-(1-\gamma)d_{\min}\alpha)^{k+1}\left\|\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{0}\right\|_{\infty}+\alpha\mathbb{E}\left[\left\|\sum^{k}_{i=0}{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\bm{\epsilon}^{\mathrm{avg}}_{i}\right\|_{\infty}\right]+\alpha\mathbb{E}\left[\sum^{k}_{i=0}\left\|{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\right\|_{\infty}\left\|{\bm{E}}_{i}\right\|_{\infty}\right]$
$\leq$ $(1-(1-\gamma)d_{\min}\alpha)^{k+1}\left\|\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{0}\right\|_{\infty}+\alpha^{\frac{1}{2}}\frac{8\sqrt{2}R_{\max}}{(1-\gamma)^{\frac{3}{2}}d_{\min}^{\frac{1}{2}}}\sqrt{\ln(2|{\mathcal{S}}||{\mathcal{A}}|)}+\alpha\mathbb{E}\left[\sum^{k}_{i=0}\left\|{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\right\|_{\infty}\left\|{\bm{E}}_{i}\right\|_{\infty}\right]$
@@
$\sum^{k}_{i=0}\left\|{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\right\|_{\infty}\left\|{\bm{E}}_{i}\right\|_{\infty}\leq$ $\gamma d_{\max}\sum^{k}_{i=0}(1-(1-\gamma)d_{\min}\alpha)^{k-i}\left(\sigma_{2}({\bm{W}})^{i}\left\|\bm{\Theta}\bar{{\bm{Q}}}_{0}\right\|_{2}+\alpha\frac{8R_{\max}}{1-\gamma}\frac{\sqrt{N|{\mathcal{S}}||{\mathcal{A}}|}}{1-\sigma_{2}({\bm{W}})}\right)$
@@
$\alpha=\tilde{{\mathcal{O}}}\left(\min\left\{\frac{(1-\gamma)^{5}d_{\min}^{3}}{R_{\max}^{2}d_{\max}^{2}}\frac{1}{{\epsilon}^{2}},\frac{(1-\gamma)^{3}d_{\min}^{2}(1-\sigma_{2}({\bm{W}}))}{R_{\max}d_{\max}^{2}\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}\frac{1}{{\epsilon}}\right\}\right).$
@@ Appendix G Appendix : Markovian observation model
where we have replaced $s_{k}^{\prime}$ in (2) with $s_{k+1}$. The overall algorithm is given in Algorithm 2 in the Appendix Section I.
@@ G.1 Analysis of optimality error under Markovian observation model
For $10\leq q\leq\tau-$ 1, let for $t\in{\mathbb{N}}$ such that $q\leq\tau t+q\leq k$:
@@
$\left\|\sum_{j=\tau}^{k}{\bm{A}}^{k-j}_{{\bm{Q}}^{*}}{\bm{w}}_{j,3}\right\|_{\infty}\leq$ $\sum^{k}_{j=\tau}(1-(1-\gamma)d_{\min}\alpha)^{k-j}\left(8\sigma_{2}({\bm{W}})^{j-\tau}\left\|\bar{{\bm{Q}}}_{0}\right\|_{2}+8\alpha\frac{8R_{\max}}{1-\gamma}\frac{\sqrt{N|{\mathcal{S}}||{\mathcal{A}}|}}{1-\sigma_{2}({\bm{W}})}+4\alpha\tau\frac{2R_{\max}}{1-\gamma}\right)$
@@
$\mathbb{E}\left[\left\|\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{k+1}\right\|_{\infty}\right]\leq$ $(1-(1-\gamma)d_{\min}\alpha)^{k-\tau+1}\mathbb{E}\left[\left\|\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{\tau}\right\|_{\infty}\right]$

### v2

@@ 4 Error Analysis : Markovian observation model
Suppose $\alpha=\tilde{{\mathcal{O}}}\left(\frac{{\epsilon}^{2}}{\ln\left(\frac{1}{{\epsilon}^{2}}\right)}\frac{(1-\gamma)^{5}d_{\min}^{3}}{t_{\mathrm{mix}}d_{\max}^{2}}\right)$. Then, the following number of samples are required for $\mathbb{E}\left[\left\|\bar{{\bm{Q}}}_{k}-\bm{1}_{N}\otimes{\bm{Q}}^{*}\right\|_{\infty}\right]\leq\epsilon$:
$\tilde{{\mathcal{O}}}\left(\max\left\{\frac{\ln^{2}\left(\frac{1}{{\epsilon}^{2}}\right)}{{\epsilon}^{2}}\frac{t_{\mathrm{mix}}d_{\max}^{2}}{(1-\gamma)^{6}d_{\min}^{4}},\frac{\ln\left(\frac{1}{{\epsilon}}\right)}{{\epsilon}}\frac{d_{\max}\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\gamma)^{4}d_{\min}^{3}(1-\sigma_{2}({\bm{W}}))}\right\}\right).$
The proof is given in Appendix Section G.3. As in the result of i.i.d. case in Corollary 3.8, we have the dependency on $\frac{1}{1-\gamma},\frac{1}{d_{\min}}$, and $\frac{1}{1-\sigma_{2}({\bm{W}})}$ with additional factor on mixing time. The known tight sample complexity result in the single-agent case is $\tilde{{\mathcal{O}}}\left(\frac{1}{(1-\gamma)^{4}d_{\min}{\epsilon}^{2}}+\frac{t_{\mathrm{mix}}}{(1-\gamma)d_{\min}}\right)$ by Li et al. [2024], and our result leaves room for improvement. Assuming a uniform sampling scheme, i.e., $d_{\min}=d_{\max}=\frac{1}{|{\mathcal{S}}||{\mathcal{A}}|}$, and $|{\mathcal{A}}_{i}|=A$ for all $i\in[N]$ and $A\geq 2$, the sample complexity becomes $\tilde{{\mathcal{O}}}\left(\max\left\{\frac{t_{\text{mix}}}{{\epsilon}^{2}}\frac{|{\mathcal{S}}|^{2}A^{2N}}{(1-\gamma)^{6}},\frac{1}{{\epsilon}}\frac{|{\mathcal{S}}|^{\frac{5}{2}}A^{\frac{5N}{2}}}{(1-\gamma)^{4}(1-\sigma_{2}({\bm{W}}))}\right\}\right)$. We note that the exponential scaling in the action space is inevitable in the tabular setting unless we consider a near-optimal solution [Qu et al., 2022]. Lastly, to verify the convergence of our algorithm, experiments are provided in Appendix Section I.
@@ 5 Discussion
Q-function Assumption Sample complexity Bound type Remarks Ours Tabular $\max\left\{\frac{t_{\text{mix}}}{\epsilon^{2}}\frac{1}{(1-\gamma)^{6}d_{\min}^{3}},\frac{1}{\epsilon}\frac{\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\sigma_{2}(\bm{W}))(1-\gamma)^{4}d_{\min}^{3}}\right\}$ Expectation -  Wang et al. [2022] Tabular $\frac{1}{(1-\gamma)^{5}d_{\min}{\epsilon}^{2}}+\frac{t_{\mathrm{mix}}}{1-\gamma}$ High probability ${\epsilon}\in\left[0,\frac{1}{1-\gamma}\right)$  Heredia et al. [2020] LFA (15) $\frac{R^{2}}{(d_{\min}-\gamma^{2}d_{\max}^{*})^{2}(1-\sigma_{2}({\bm{W}}))}$ Expectation Averaged squared error Continuous state space $R$ is projection radius  Zeng et al. [2022b] LFA (16) $\frac{1}{\kappa^{2}(1-\gamma)^{2}(1-\sigma_{2}({\bm{W}}))}$ Expectation -
Table 1: LFA stands for linear function approximation.
@@
for some $\kappa>0$. We have provided examples where the above conditions in (15) and (16) are not met even in the tabular case in Appendix Section H.
Overall, the assumptions used in Heredia et al. [2020], Zeng et al. [2022b] allows the analysis to follow similar lines to that of convex optimization literature. To the best of our knowledge, there is no existing literature that demonstrates how to extend convex optimization analysis, or an analogous approach, to the analysis of Q-learning under the tabular setup. This gap in the literature makes the analysis challenging and is the primary reason we rely on switched system analysis. Due to different settings, their sample complexity is not directly comparable with ours.
Wang et al. [2022] proposed a distributed Q-learning algorithm in the tabular setting, which is motivated from the adapt-then-combine algorithm, whereas our algorithm considers combine-and-adapt scheme [Chen and Sayed, 2012] in the distributed optimization literature. The work presents a sharper bound on the sample complexity $\frac{1}{(1-\gamma)^{5}d_{\min}\epsilon^{2}}$ compared to ours $\frac{1}{(1-\gamma)^{6}d_{\min}^{4}\epsilon^{2}}$ Nonetheless, the algorithm proposed by Wang et al. [2022] requires two steps for a single update, whereas in our paper, we focus on a one-step algorithm that is algorithmically simpler and more efficient. The communication process must occur after the update scheme, whereas in our model, the TD-error computation can be performed concurrently with the communication process. Specifically, we analyze the traditional and widely adopted QD-learning algorithm proposed in Kar et al. [2013], for which a finite-time error analysis for the original form has been lacking in the literature. Additionally, we enhance the efficiency of QD-learning by employing a constant step-size, as opposed to the two-time-scale decaying step-size used in traditional QD-learning. This modification can significantly improve the initial convergence speed empirically.
@@ 6 Conclusion
In this paper, we have studied distributed version of Q-learning algorithm. We provided a sample complexity result of $\tilde{\mathcal{O}}\left(\max\left\{\frac{1}{\epsilon^{2}}\frac{1}{(1-\gamma)^{6}d_{\min}^{4}},\frac{1}{\epsilon}\frac{\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\sigma_{2}(\bm{W}))(1-\gamma)^{4}d_{\min}^{3}}\right\}\right)$, which appears to be the first non-asymptotic result for tabular Q-learning. Future work would include improving the dependency on $\frac{1}{1-\gamma}$ and $d_{\min}$ to match the known tightest sample complexity bound of single-agent Q-learning [Li et al., 2020]. Furthermore, to resolve the scalability issue, two promising approaches would be adopting a mean-field approach or exploring convergence to sub-optimal point.
7 Acknowledgments
The work was supported by the Institute of Information Communications Technology Planning Eval- uation (IITP) funded by the Korea government under Grant 2022-0-00469.
Appendix A Appendix : Notations and Organization
Notations:
[…]
Organization:
The paper is organized as follows: Section 2 provides background for the MARL setting. Section 3 provides result under i.i.d. observation model and sketch of the proof. The result for Markovian observation model is provided in Section 4.
@@ Appendix B Appendix : Constructing Doubly Stochastic Matrix
One can formulate a semi-definite program to construct a doubly stochastic matrix [Xiao and Boyd, 2004]. It finds the doubly stochastic matrix with minimum possible $\sigma_{2}(W)$ but it requires a centralized controller to solve such system, and distributed the computed the result of each agents. Another choice is to use Sinkhorn-Knopp algorithm [Knight, 2008]. However, it also requires a centralized computation scheme. Moreover, to our best knowledge, we are not aware of bound on the $\sigma_{2}(W)$ of the output of Sinkhorn-Knopp algorithm.
@@ Appendix D Appendix : Omitted Proofs
To prove the inequality in (17), we first prove that 111 is the unique largest eigenvalue of ${\bm{W}}$. Noting that $\bm{1}_{N}$ is an eigenvector of ${\bm{W}}$ with eigenvalue of 111, and $\rho({\bm{W}})\leq||{\bm{W}}||_{\infty}=1$ where $\rho(\cdot)$ is the spectral radius of a matrix, the largest eigenvalue of ${\bm{W}}$ should be one. This implies that $\sigma_{2}({\bm{W}})<1$. The multiplicity of the eigenvalue 111 is one, which follows from the fact that ${\bm{W}}^{k}$ is a non-negative and irreducible matrix and that the largest eigenvalue of a non-negative and irreducible matrix is unique Pillai et al. [2005] from Perron-Frobenius theorem. Note that ${\bm{W}}^{k}$ is a non-negative and irreducible matrix due to the fact that the graph ${\mathcal{G}}$ is connected.
@@ E.2 Construction of upper comparison system
$\mathbb{E}\left[{\bm{S}}_{0}\right]=$ $\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}({\bm{e}}_{s_{0}}\otimes{\bm{e}}_{{\bm{a}}_{0}})(r^{i}_{1}+{\bm{e}}_{s^{\prime}_{0}}^{\top}\gamma{\bm{\Pi}}^{{\bm{Q}}^{i}_{0}}{\bm{Q}}^{i}_{0}-({\bm{e}}_{s_{0}}\otimes{\bm{e}}_{{\bm{a}}_{0}})^{\top}{\bm{Q}}^{i}_{0})\right]$
$-\mathbb{E}\left[\frac{1}{N}\sum^{N}_{i=1}{\bm{D}}({\bm{R}}^{i}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}^{i}_{0}}{\bm{Q}}^{i}_{k}-{\bm{Q}}^{i}_{0})\right]$
@@
$\leq$ $(1-(1-\gamma)d_{\min}\alpha)^{k+1}\left\|\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{0}\right\|_{\infty}+\alpha\mathbb{E}\left[\left\|\sum^{k}_{i=0}{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\bm{\epsilon}^{\mathrm{avg}}_{i}\right\|_{\infty}\right]$
$+\alpha\mathbb{E}\left[\sum^{k}_{i=0}\left\|{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\right\|_{\infty}\left\|{\bm{E}}_{i}\right\|_{\infty}\right]$
$+\alpha\mathbb{E}\left[\sum^{k}_{i=0}\left\|{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\right\|_{\infty}\left\|{\bm{E}}_{i}\right\|_{\infty}\right]$
@@
$s\sum^{k}_{i=0}\left\|{\bm{A}}^{k-i}_{{\bm{Q}}^{*}}\right\|_{\infty}\left\|{\bm{E}}_{i}\right\|_{\infty}$
$\leq$ $\gamma d_{\max}\sum^{k}_{i=0}(1-(1-\gamma)d_{\min}\alpha)^{k-i}\left(\sigma_{2}({\bm{W}})^{i}\left\|\bm{\Theta}\bar{{\bm{Q}}}_{0}\right\|_{2}+\alpha\frac{8R_{\max}}{1-\gamma}\frac{\sqrt{N|{\mathcal{S}}||{\mathcal{A}}|}}{1-\sigma_{2}({\bm{W}})}\right)$
@@
$\alpha=\tilde{{\mathcal{O}}}\left(\min\left\{\frac{(1-\gamma)^{5}d_{\min}^{3}}{R_{\max}^{2}d_{\max}^{2}}{\epsilon}^{2},\frac{(1-\gamma)^{3}d_{\min}^{2}(1-\sigma_{2}({\bm{W}}))}{R_{\max}d_{\max}^{2}\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{\epsilon}\right\}\right).$
@@ Appendix G Appendix : Markovian observation model
where we have replaced $s_{k}^{\prime}$ in (2) with $s_{k+1}$. The overall algorithm is given in Algorithm 2 in the Appendix Section J.
@@ G.1 Analysis of optimality error under Markovian observation model
For $0\leq q\leq\tau-1$, let for $t\in{\mathbb{N}}$ such that $q\leq\tau t+q\leq k$:
@@
$\left\|\sum_{j=\tau}^{k}{\bm{A}}^{k-j}_{{\bm{Q}}^{*}}{\bm{w}}_{j,3}\right\|_{\infty}$
$\leq$ $\sum^{k}_{j=\tau}(1-(1-\gamma)d_{\min}\alpha)^{k-j}\left(8\sigma_{2}({\bm{W}})^{j-\tau}\left\|\bar{{\bm{Q}}}_{0}\right\|_{2}+8\alpha\frac{8R_{\max}}{1-\gamma}\frac{\sqrt{N|{\mathcal{S}}||{\mathcal{A}}|}}{1-\sigma_{2}({\bm{W}})}+4\alpha\tau\frac{2R_{\max}}{1-\gamma}\right)$
@@
$\leq$ $(1-(1-\gamma)d_{\min}\alpha)^{k-\tau+1}\mathbb{E}\left[\left\|\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{\tau}\right\|_{\infty}\right]$